carbon dioxide gas enters a pipe at 3MPa and 500K at a rate of 2kg/s. CO2 is cooled at constant pressure as it flows in the pipe and the temperature CO2 drops to 450K at the exit. determine the volume flow rate and the density of CO2 at the inlet and the volume flow rate at the exit of the pipe using (a) the ideal-gas equation and (b) the generalized compressibility chart.

Grade:Upto college level

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Pi = 3MPa,Ti=500K,mi=2kg/s

PV = nRT

=> volume flow rate at inlet = miRT/PM

=> 2*8.314*10^3*500/44*3*10^6

=> 0.06295 m^3/sec

density at inlet is = PM/RT

=> 3*10^6*44/8.314*500

=> 31.75 kg/m^3

density at exit = 3*10^3*44/8.314*450 = 35.28 kg/m^3